OpenStudy (katherinesmith):

true or false: The following is the graph of the parent function y = sqrt(x)

OpenStudy (katherinesmith):

\[y = \sqrt{x}\]

hero (hero):

What do you think? Is it true or false?

OpenStudy (katherinesmith):

false?

hero (hero):

Yes because sqrt(0) = 0, not -2.

OpenStudy (katherinesmith):

thank youuuuu

OpenStudy (katherinesmith):

what about this one: the following is the graph of \[y = \sqrt[3]{x}\]

OpenStudy (katherinesmith):

@Hero

hero (hero):

Basically, what you have to do is find one instance where the function isn't true. And if it isn't true for that one instance, then it is false.

hero (hero):

for this particular function, what is \(\sqrt[3]{0}\) ?

OpenStudy (katherinesmith):

0?

hero (hero):

Yes. Does that match the graph?

OpenStudy (katherinesmith):

no

hero (hero):

So that means it must be false

OpenStudy (katherinesmith):

yay thank you